OpenStudy (anonymous):

Jake says that the function is defined at x = –1, x = 3, and x = 4. Yoe says that the function is undefined at those x values. Who is incorrect? Justify your reasoning.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

f(x)=(x-1)(x+2)(x+4)/(x+1)(x-2)(x-4)

OpenStudy (anonymous):

do you understand the function I wrote or should I draw i?

OpenStudy (michele_laino):

from the text of your problem I can write this expression for function f(x): \[\Large f\left( x \right) = \frac{{\left( {x - 1} \right)\left( {x + 2} \right)\left( {x + 4} \right)}}{{\left( {x + 1} \right)\left( {x - 2} \right)\left( {x - 4} \right)}}\]

OpenStudy (michele_laino):

now we can not divide by zero, so we have to exclude those values of x, such that: \[\Large \begin{gathered} x + 1 = 0 \hfill \\ x - 2 = 0 \hfill \\ x - 4 = 0 \hfill \\ \end{gathered} \] in other words we have to be certain that the denominator is not equal to zero

OpenStudy (anonymous):

ok, i see

OpenStudy (michele_laino):

for example if I solve the first equation, I get: x=-1 so I have to exclude that value

OpenStudy (michele_laino):

If I solve the second equation, I get: x=2 again, I have to exclude the value x=2

OpenStudy (michele_laino):

similarly for third equation: x-4=0 which gives: x=4, and i have to exclude that value

OpenStudy (anonymous):

ok:) I understand

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

So the answer would be? @Michele_Laino

OpenStudy (michele_laino):

Edward is right!

OpenStudy (anonymous):

ok:) Thanks!!!

OpenStudy (michele_laino):

:)